ok here is a problem that is driving me nuts i can't understand what to do. the question says " for the given functions f and g, find the indicated composition." and here is the problem f(x)=x^2+2x+3, g(x)=x^2+2x+2 (f of g)(-5)
Answer by venugopalramana(3286)

ok here is a problem that is driving me nuts i can't understand what to do. the question says " for the given functions f and g, find the indicated composition." and here is the problem f(x)=x^2+2x+3
g(x)=x^2+2x+2
f of g(-5)=f{g(-5)}
g(-5)=(-5)^2+2(-5)+2=25-10+2=17
f{g(-5)}=f(17)=17^2+2*17+3=289+34+3=326